If Greg buys gas at $2.99 per gallon. And uses 60 gallons of gas per week, how much would he pay for gas per year?
How would write this in equation form? Then how would I graph the equation for this?
PLEASE HELP?

Answer by checkley75(3666)

Cost per gallon*number of gallons per week*number of weeks.
2.99*60*52=$9,328.90 cost of 1 years worth of gas.
Sorry don't have graphing capability for these type of problems.
--------------------------------------------------------------------
IF YOU WERE TO GRAPH THIS PROBLEM THE GRAPH WOULD BE A STRAIGHT LINE ORIGINATING AT (0,0). THE X AXIS COULD BE THE WEEKS & THE Y AXIS THE COST.
FOR THE FIRST WEEK THE COST WOULD BE (2.99*60)=179.4.
FOR THE SECOND WEEK THE COST WOULD BE 2*179.4=358.8.
THE SLOPE WOULD DEPEND ON THE PRICE SCALE.
HOPE THIS HELPS.